# Beyond Functional Correctness: Investigating Coding Style Inconsistencies in Large Language Models

### v1

@@ Abstract.
Large language models (LLMs) have brought a paradigm shift to the field of code generation, offering the potential to enhance the software development process. However, previous research mainly focuses on the accuracy of code generation, while coding style differences between LLMs and human developers remain under-explored. In this paper, we empirically analyze the differences in coding style between the code generated by mainstream Code LLMs and the code written by human developers, and summarize coding style inconsistency taxonomy. Specifically, we first summarize the types of coding style inconsistencies by manually analyzing a large number of generation results. We then compare the code generated by Code LLMs with the code written by human programmers in terms of readability, conciseness, and robustness. The results reveal that LLMs and developers have different coding styles. Additionally, we study the possible causes of these inconsistencies and provide some solutions to alleviate the problem.
@@
copyright: acmlicensedjournalyear: 2024
@@ 1. Introduction
Code*Corresponding author. generation is to automatically generate code snippets that align with given requirements, which plays a vital role in the software engineering domain (OpenAI, 2021; Zhang et al., 2023e; Radford et al., 2018; Dong et al., 2023; Jiang et al., 2023b; Zan et al., 2023b, c; Olausson et al., 2023; Le et al., 2022; Muennighoff et al., 2023; Zan et al., 2023a; Huang et al., 2023; Jiang et al., 2021; Agrawal et al., 2023; Yu et al., 2023; Roziere et al., 2023a; Luo et al., 2023; Phind, 2023; Li et al., 2023a, e; Liu et al., 2023a; Zhang et al., 2023c; Sun et al., 2023a; Zheng et al., 2023b; Zhang et al., 2023a; Jain et al., 2023; Bairi et al., 2023; Zhang et al., 2023b; Ni et al., 2023; Chen et al., 2023; Li et al., 2023b; Yadav et al., 2023; Li et al., 2023d, f; Jiang et al., 2023a; Zhang et al., 2023d; Zhu et al., 2023; Wang et al., 2023b; Li et al., 2023c; Wang and Li, 2021; Shi et al., 2023a, b; Mu et al., 2023; Zheng et al., 2023c; Dong et al., 2023; Jiang et al., 2023b; Zhang et al., 2023b; Hong et al., 2024; Qian et al., 2023). Recently, the advent of large language models for code (Code LLMs) (Chen et al., 2021; Nijkamp et al., 2022; Roziere et al., 2023b; Li et al., 2023a; Guo et al., 2024a), such as CodeLlama (Roziere et al., 2023a), StarCoder (Li et al., 2023a), Codex (OpenAI, 2021), has greatly advanced the performance of code generation. These models have demonstrated remarkable capabilities in code generation, thereby significantly improving software development efficiency. However, previous studies mainly focus on improving the accuracy of LLM-based code generation, another important aspect, coding style of Code LLMs, remains under-explored. Understanding the coding style differences between Code LLMs and human developers is crucial, as the coding style can affect code readability, maintainability, and overall software quality.
@@
There are several previous works related to coding style (Oman and Cook, 1990; Parr and Vinju, 2016; Markovtsev et al., 2019; Mi et al., 2016; Chen and Abedjan, 2023). Oman et al. (Oman and Cook, 1990) proposed a programming style taxonomy, but this taxonomy may be outdated in the era of LLMs. CODEBUFF (Parr and Vinju, 2016), an automatic code formatter, and STYLE-ANALYZER (Markovtsev et al., 2019), which repairs code formatting inconsistencies, focus exclusively on code formatting style. Mi et al. (Mi et al., 2016) expanded the scope by using hierarchical agglomerative clustering to measure stylistic inconsistency, considering not only formatting but also stylistic metrics related to code readability and features specific to the C/C++ programming languages. More recently, DUETCS (Chen and Abedjan, 2023) was proposed for coding style transfer. This work considers a broader range of coding style features, categorizing them into text style (formatting and naming conventions) and structure style (code blocks ordering and preferences of control flow statements). These works provide a preliminary foundation and inspiration for studying coding styles. However, several gaps remain. Firstly, the classification and definition of coding styles remain insufficiently detailed and comprehensive. Additionally, no existing research has analyzed the differences in coding style between Code LLMs and human developers. Furthermore, there has been no comparative analysis of the coding styles among different Code LLMs.
@@
In this paper, we aim to fill these gaps by conducting the first empirical study to examine inconsistencies in coding styles between code generated by mainstream Code LLMs and code written by human developers. ① Firstly, we conduct extensive manual analysis to categorize various types of coding style inconsistencies. Specifically, we compare the code generation results of four mainstream Code LLMs (CodeLlama-7B (Roziere et al., 2023a), StarCoder2-7B (Li et al., 2023a), DeepSeekCoder-1.3B (Guo et al., 2024b), and DeepSeekCoder-6.7B (Guo et al., 2024b)) with ground truth on the CoderEval (Yu et al., 2024) benchmark. We annotate the results and perform open coding to obtain a comprehensive taxonomy of coding style inconsistencies. [^1] In this paper, we use “coding style inconsistency”, “style inconsistency”, and “inconsistency” interchangeably. ② Secondly, we analyze the distribution of the inconsistencies, including the inconsistency ratio, frequency, and differences for different Code LLMs. ③ Thirdly, we investigate which coding style is better by comparing the generated code against human-written code across several dimensions, including readability, conciseness, and robustness. ④ Finally, we experiment on several prompting strategies to explore methods to improve the coding style of Code LLMs.
@@
Through extensive experiments and evaluation, we have obtained the following results on coding style inconsistencies of Code LLMs generated code. 1 We propose the first coding style inconsistency taxonomy of Code LLM-based code generation. The taxonomy contains 24 inconsistency types that cover all inconsistency cases in the studied LLMs. We further categorize the 24 inconsistency types into five dimensions, i.e., Formatting Inconsistency, Semantic Inconsistency, Expression/Statement Inconsistency, Control Follow Inconsistency, and Fault Tolerance Inconsistency. 2 Analysis results indicate that there are obvious coding style inconsistencies between human and all studied Code LLMs, especially in statements/expressions and formatting dimensions. In addition, coding styles of Code LLMs themselves are generally similar, although there are some differences in the formatting dimension. 3 Overall, code generated by Code LLMs is comparable to or even better than human-written code in terms of readability, conciseness, and robustness. 4 We find that certain types of prompts can slightly improve the readability and robustness of generated code. But there is a trade-off between readability and conciseness, indicating that while prompt engineering can help, it is not sufficient to fully address issues related to coding style and sometimes it may even decrease code generation accuracy. 5 Through case studies, we carefully analyze several common scenarios of coding style inconsistencies in Code LLMs, and find that they tend to use deprecated APIs, are unfamiliar with basic Python functions, and rarely use advanced syntax features, resulting in less concise and less efficient code.
@@
  • We provide a comprehensive taxonomy of coding style inconsistencies between Code LLMs and human developers.
  • We conduct extensive analysis to reveal the coding style inconsistencies of mainstream Code LLMs, leading to a deeper understanding of LLM-based code generation.
  • We propose practical solutions to improve coding style discrepancies, paving the way for a more harmonious integration of LLMs and coding practices.
@@ 2.1. LLM-based Code Generation
Code LLMs, such as StarCoder (Li et al., 2023a), CodeLlama (Roziere et al., 2023a), and DeepSeek-Coder (Guo et al., 2024a), are specifically optimized for code-centric tasks (Zheng et al., 2023b, a), leveraging massive code-specific corpora and specialized training instructions. In recent years, some works have studied the application of Code LLMs in fields such as vulnerability detection (Cheshkov et al., 2023; Wang et al., 2024b; Tamberg and Bahsi, 2024; Yang et al., 2024; Yusuf and Jiang, 2024), commit message generation (Lopes et al., 2024; Mandli et al., 2024; Tao et al., 2024; Zhang et al., 2024b), unit test generation (Siddiq et al., 2023; Xie et al., 2023; Yuan et al., 2023b; Schäfer et al., 2023), code search (Kondo et al., 2024; Wang et al., 2023a; Hu et al., 2024; Li et al., 2024d; Gong et al., 2024), code summarization (Sun et al., 2023b; Ahmed et al., 2024; Su and McMillan, 2024; Li et al., 2024c; Haldar and Hockenmaier, 2024; Wang et al., 2024a) and code generation (Liu et al., 2023c; Yuan et al., 2023a; Liu et al., 2024; Huang et al., 2024; Sun et al., 2024; Zhang et al., 2024a; Zhu et al., 2024; Tipirneni et al., 2024; Ugare et al., 2024; Zhuo et al., 2024), etc.
To understand the code generation performance of Code LLMs, some high-quality code generation benchmarks have been proposed in recent years. For example, HumanEval (Chen et al., 2021), MBPP (Austin et al., 2021), ClassEval (Du et al., 2023), covering different scenarios such as repository-level code generation (Zhang et al., 2023c; Phan et al., 2024; Li et al., 2024b, a) and class-level code generation tasks (Du et al., 2023). While most studies are primarily concerned with improving the functional correctness of code generated by models, using metrics like passk (Chen et al., 2021), recent research has begun to explore other attributes of code generated by Code LLMs. For instance, methods have been proposed to enhance the robustness of Code LLMs (Chen et al., 2022; Zhang et al., 2022), and attention has been given to the security aspects of Code LLMs in code generation tasks, investigating potential vulnerabilities and risks (Oh et al., 2023; Cotroneo et al., 2023).
In contrast to previous works, our investigation is on the code style of Code LLMs. We conduct the first study to compare the code style of several mainstream Code LLMs with code written by human programmers. Additionally, we compare the code styles among different mainstream Code LLMs. This analysis provides insights into the strengths and weaknesses of Code LLMs in terms of coding style, shedding light on potential areas for improvement and future research directions.
@@ 2.2. Coding Style
In previous work  (Oman and Cook, 1990), Oman et al. established a programming style taxonomy, a cornerstone for developing programming style guidelines and analyzers. Recent strides in coding style research include innovations like CODEBUFF (Parr and Vinju, 2016), an automatic code formatter that leverages machine learning to understand and apply code formatting styles. Similarly, STYLE-ANALYZER (Markovtsev et al., 2019) addresses code formatting inconsistencies using a decision tree forest model. However, both CODEBUFF and STYLE-ANALYZER focus solely on formatting style.
Mi et al. (Mi et al., 2016) employed hierarchical agglomerative clustering to gauge code style inconsistencies, focusing on C/C++ languages. In a recent study (Chen and Abedjan, 2023), DUETCS extracted comprehensive code style features from target code examples, covering text and structure style elements. DUETCS utilizes a Siamese feature network to transform source code style into that of target examples while preserving semantic integrity.
Unlike previous studies, our work represents the first empirical examination of coding style inconsistencies between code generated by Code LLMs and code written by human programmers. Drawing on established coding style categories and definitions from prior literature, we conducted open coding on samples generated by several mainstream Code LLMs. This process yielded a coding style inconsistency taxonomy comprising five dimensions and 24 distinct inconsistency types. In comparison to prior efforts, our proposed terminology of code style inconsistencies is more comprehensive and detailed, extending beyond traditional considerations of text style and structure. Furthermore, our study lays the groundwork for future research on the coding style of Code LLMs, offering valuable insights and avenues for further exploration in this field.
@@ 3. Experimental Setup
In this section, we introduce the experimental setup, including the Code LLM selection, dataset description, and implementation details.
3.1. Code LLM Selection
We select four mainstream and representative open-sourced Code LLMs that have demonstrated strong performance in the code generation task, namely CodeLlama-7B, StarCoder2-7B, DeepSeekCoder-1.3B, and DeepSeekCoder-6.7B. Due to the constraints in computing resources, we exclude larger models with more than 7 billion parameters. The models we selected are all base models without instruction-tuning, which is particularly suitable for our code completion scenario, wherein the task is to complete the code based on the given context. For the four selected Code LLMs, we directly obtain and run their released versions from their official repositories, following the provided documentation. The same settings are being used for all LLMs.
@@ 3.2. Benchmark Selection
Our experiments are conducted on CoderEval (Yu et al., 2024), which is a benchmark used to evaluate code generation performance on pragmatic code generation tasks, i.e., code generation with repository context. It consists of 230 Python and 230 Java tasks from real-world open-source projects. Each task contains a function signature [^2] We use “method” and “function” interchangeably in this paper., a task description, a solution code as the ground truth, and several unit tests to assess the functional correctness of the generated code. The objective of each task is to complete the code specified by the function signature, guided by the provided task description, and ensure that it passes the associated unit tests. In this study, we focus on Python tasks due to Python’s popularity (Srinath, 2017) and its alignment with previous code generation work (Bafatakis et al., 2019).
3.3. Experimental Details
We configure all the Code LLMs to use the same hyperparameter settings. We adopt a random sampling strategy and set the maximum context length to 1024 and the temperature to 0.6. Each model generates one output sequence at a time. The settings for the maximum window length and temperature are based on the experimental setup as in CoderEval (Yu et al., 2024). All experiments are conducted on a machine with 216 GB main memory and a Tesla A100 GPU with 80GB memory.

### v2

@@ Abstract.
Large language models (LLMs) have brought a paradigm shift to the field of code generation, offering the potential to enhance the software development process. However, previous research mainly focuses on the accuracy of code generation, while coding style differences between LLMs and human developers remain under-explored. In this paper, we empirically analyze the differences in coding style between the code generated by mainstream LLMs and the code written by human developers, and summarize coding style inconsistency taxonomy. Specifically, we first summarize the types of coding style inconsistencies by manually analyzing a large number of generation results. We then compare the code generated by LLMs with the code written by human programmers in terms of readability, conciseness, and robustness. The results reveal that LLMs and developers exhibit differences in coding style. Additionally, we study the possible causes of these inconsistencies and provide some solutions to alleviate the problem.
@@
copyright: rightsretainedjournal: PACMSEjournalyear: 2025journalvolume: 2journalnumber: FSEarticle: FSE032publicationmonth: 7doi: 10.1145/3715749ccs: Software and its engineering Software evolutionccs: Software and its engineering Automatic programming
@@ 1. Introduction
Code generation is to automatically generate code snippets that align with given requirements, which plays a vital role in the software engineering domain (Dong et al., 2023; Zan et al., 2023b; Olausson et al., 2023; Le et al., 2022; Zan et al., 2023a; Huang et al., 2023; Jiang et al., 2021; Li et al., 2023e; Bairi et al., 2023; Chen et al., 2023; Li et al., 2023b; Yadav et al., 2023; Li et al., 2023d, f; Jiang et al., 2023a; Zhang et al., 2023c; Shi et al., 2023; Mu et al., 2023; Zheng et al., 2023b; Jiang et al., 2023b; Zhang et al., 2023a). Recently, the advent of large language models  (Nijkamp et al., 2022; Li et al., 2023a; Guo et al., 2024b), such as CodeLlama (Roziere et al., 2023), StarCoder (Li et al., 2023a), Codex (OpenAI, 2021) and GPT-4 (Achiam et al., 2023), has greatly advanced the performance of code generation. These models have demonstrated remarkable capabilities in code generation, significantly enhancing software development efficiency (Zheng et al., 2023c). However, while previous studies primarily focus on enhancing the accuracy of LLM-based code generation, an equally important aspect—coding style—remains under-explored. Understanding the differences in coding style between LLMs and human developers is essential, as it directly impacts code readability, maintainability, and overall software quality (Mi et al., 2016).
@@
There are several previous works related to coding style (Oman and Cook, 1990; Parr and Vinju, 2016; Markovtsev et al., 2019; Mi et al., 2016; Chen and Abedjan, 2023). Oman et al. (Oman and Cook, 1990) proposed a programming style taxonomy, including typographic style, control structure style, and information structure style. However, some of these guidelines are outdated, such as the use of Goto statements, which are now rarely used in modern programming. To improve code formatting style, tools like CODEBUFF (Parr and Vinju, 2016), an automatic code formatter, and STYLE-ANALYZER (Markovtsev et al., 2019), which fixes formatting inconsistencies, have been developed. Mi et al. (Mi et al., 2016) expanded the scope by using hierarchical agglomerative clustering to measure stylistic inconsistency, considering not only formatting but also stylistic metrics related to code readability and features specific to the C/C++ programming languages. More recently, DUETCS (Chen and Abedjan, 2023) was proposed for coding style transfer. This work considers a broader range of coding style features, categorizing them into text style (formatting and naming conventions) and structure style (code blocks ordering and preferences of control flow statements). These works provide a preliminary foundation and inspiration for examining coding styles. In our work, we adopt the definition of code style in previous research (Mi et al., 2016; Berry and Meekings, 1985; Yasir and Kabir, 2022; Oman and Cook, 1990): Code style can be understood as a personal habit in writing source code, reflecting individual preferences in aspects like physical layout, algorithms, etc. However, there remain several gaps. (i) Existing work offers a generalized definition of code style but lacks detailed and comprehensive classification. (ii) No existing studies have yet analyzed the differences in coding style between mainstreams LLMs and human developers. (iii) A comparative analysis of coding styles across different LLMs is still missing.
@@
In this paper, our aim is to fill these gaps. ① Firstly, we conduct extensive manual analysis to categorize various types of coding style inconsistencies. We compare the code generation results of five mainstream LLMs, including four Code LLMs and a state-of-the-art general-purpose LLM (GPT-4 (Achiam et al., 2023)), against the ground truths of CoderEval (Yu et al., 2024) benchmark. We annotate the results and perform open coding to develop a comprehensive taxonomy of coding style inconsistencies, achieving a detailed classification of code style, which addresses (i) and (ii).  [^1] In this paper, we use “coding style inconsistency”, “style inconsistency”, and “inconsistency” interchangeably. ② Secondly, we analyze the distribution of inconsistencies, including ratios, frequencies, and differences across the five LLMs, addressing (iii). ③ Thirdly, we compare the generated code with human-written code based on readability, conciseness, and robustness. ④ Finally, we experiment on several prompting strategies to explore methods to improve the coding style of LLMs.
@@
Through extensive experiments and evaluation, we have obtained the following results: 1 We propose the first taxonomy of coding style inconsistencies in LLM-generated code. The taxonomy contains 24 inconsistency types, categorized into five dimensions, i.e., Formatting Inconsistency, Semantic Inconsistency, Expression/Statement Inconsistency, Control Follow Inconsistency, and Fault Tolerance Inconsistency. 2 Our analysis reveals significant style inconsistencies between human-written code and all studied LLMs, particularly in statements/expressions and formatting. Coding styles among the LLMs themselves are generally similar, with some variation in formatting. 3 Overall, LLM-generated code is comparable to human-written code in terms of readability, conciseness, and robustness. 4 Certain prompt strategies can marginally improve the readability and robustness of generated code, but there is a trade-off with conciseness. This suggests that while prompt engineering can help, it may not fully resolve coding style issues and can sometimes reduce code accuracy. 5 Our case studies show that LLM-generated code often exhibits poor formatting, lacks familiarity with basic Python functions, and underutilizes advanced syntax features, making the code less readable, concise, and efficient.
@@
  • We provide a comprehensive taxonomy of coding style inconsistencies between LLMs and human developers. Compared to previous work, our proposed taxonomy achieves a more detailed classification of coding styles.
  • We conduct extensive analysis to reveal the coding style inconsistencies of several mainstream LLMs, leading to a deeper understanding of LLM-based code generation.
  • We propose some practical solutions to improve the inconsistencies of the coding style, paving the way for a more harmonious integration of LLM and coding practices.
@@ 2.1. LLM-based Code Generation
Code LLMs, such as StarCoder (Li et al., 2023a), CodeLlama (Roziere et al., 2023), and DeepSeek-Coder (Guo et al., 2024b), are specifically optimized for code-centric tasks (Zheng et al., 2023c, a), leveraging massive code-specific corpora and specialized training instructions. The most state-of-the-art general-purpose large model, GPT-4, due to its significantly larger number of parameters compared to other Code LLMs, demonstrates outstanding performance on code-centric tasks. In recent years, some works have studied the application of LLMs in fields such as vulnerability detection (Cheshkov et al., 2023; Wang et al., 2024b; Tamberg and Bahsi, 2024; Yang et al., 2024; Yusuf and Jiang, 2024), commit message generation (Lopes et al., 2024; Mandli et al., 2024; Tao et al., 2024; Zhang et al., 2024), unit test generation (Siddiq et al., 2023; Xie et al., 2023; Yuan et al., 2023b; Schäfer et al., 2023), code search (Kondo et al., 2024; Wang et al., 2023a; Hu et al., 2024; Li et al., 2024e; Gong et al., 2024), code summarization (Sun et al., 2023b; Ahmed et al., 2024; Su and McMillan, 2024; Li et al., 2024d; Haldar and Hockenmaier, 2024; Wang et al., 2024a) and code generation (Liu et al., 2023b; Yuan et al., 2023a; Liu et al., 2024; Huang et al., 2024; Zhu et al., 2024; Tipirneni et al., 2024; Ugare et al., 2024; Sun et al., 2023a; Jain et al., 2023; Ni et al., 2023; Li et al., 2023c; Wang et al., 2023b; Guo et al., 2024a; Wang et al., 2024c; Li et al., 2024c), etc.
To understand the code generation performance of LLMs, some high-quality code generation benchmarks have been proposed in recent years. For example, HumanEval (Chen et al., 2021), MBPP (Austin et al., 2021) and ClassEval (Du et al., 2023), covering different scenarios such as repository-level code generation (Zhang et al., 2023b; Phan et al., 2024; Li et al., 2024b, a) and class-level code generation tasks (Du et al., 2023). While most studies are primarily concerned with improving the functional correctness of code generated by models, using metrics like passk (Chen et al., 2021), recent research has begun to explore other attributes of code generated by LLMs. For instance, methods have been proposed to enhance the robustness of LLMs (Chen et al., 2022; Zhang et al., 2022), and attention has been given to the security aspects of LLMs in code generation tasks, investigating potential vulnerabilities and risks (Oh et al., 2023; Cotroneo et al., 2023).
In contrast to previous works, our investigation is on the code style of LLMs. We conduct the first study to compare the code style of several mainstream LLMs with code written by human programmers. Additionally, we compare the code styles among different mainstream LLMs. This analysis provides insights into the strengths and weaknesses of LLMs in terms of coding style, shedding light on potential areas for improvement and future research directions.
@@ 2.2. Coding Style
In previous work (Oman and Cook, 1990), Oman et al. proposed a programming style taxonomy, encompassing typographic style, control structure style, and information structure style, which laid the foundation for the development of programming style guidelines and analyzers. However, certain rules in this taxonomy may now be considered outdated and may not fully reflect modern programming practices and conventions (For instance, the example rules mentioning the use of Goto statements are no longer widely used in contemporary programming). Recent strides in coding style research include innovations like CODEBUFF (Parr and Vinju, 2016), an automatic code formatter that leverages machine learning to understand and apply code formatting styles. Similarly, STYLE-ANALYZER (Markovtsev et al., 2019) addresses code formatting inconsistencies using a decision tree forest model. However, both CODEBUFF and STYLE-ANALYZER focus solely on formatting style. Mi et al. (Mi et al., 2016) employed hierarchical agglomerative clustering to gauge code style inconsistencies, focusing on C/C++ languages. In a recent study (Chen and Abedjan, 2023), DUETCS extracted comprehensive code style features from target code examples, covering text and structure style elements. DUETCS utilizes a Siamese feature network to transform source code style into that of target examples while preserving semantic integrity.
In our study, we adopt the definition of code style provided by earlier works (Mi et al., 2016; Berry and Meekings, 1985; Yasir and Kabir, 2022; Oman and Cook, 1990): Code style can be understood as a personal writing habit in source code, reflecting individual preferences in aspects like physical layout, algorithms, and more. While previous research offers a generalized definition of code style, the classification of code style lacks sufficient detail and comprehensiveness. To address this gap, we conducted open coding on code samples generated by several mainstream LLMs. This process result a coding style inconsistency taxonomy comprising 5 dimensions and 24 distinct inconsistency types, offering a comprehensive classification for code style. In comparison to prior efforts, our proposed taxonomy of code style inconsistencies is more comprehensive and detailed, extending beyond traditional focus on text style and structure style. Moreover, our study lays the groundwork for future research on the coding style of LLMs, offering valuable insights and avenues for further exploration in this domain.
@@ 3. Experimental Setup
In this section, we introduce the experimental setup, including the LLM selection, benchmark selection, and implementation details.
3.1. LLM Selection
We select five LLMs for our experiments, including four well-known Code LLMs (CodeLlama-7B (Roziere et al., 2023), StarCoder2-7B (Li et al., 2023a), DeepSeekCoder-1.3B (Guo et al., 2024c), and DeepSeekCoder-6.7B (Guo et al., 2024c)) and the state-of-the-art general-purpose model, GPT-4 (Achiam et al., 2023). The rationale behind choosing these four Code LLMs is their demonstrated strong performance in code generation tasks. As for GPT-4, it is chosen due to its status as the most effective and parameter-intensive model. The four selected code LLMs are base models without instruction-tuning, making them particularly well-suited for our code generation scenario, where the objective is to generate code based on the provided context.
@@ 3.2. Benchmark Selection
Our experiments are conducted on CoderEval (Yu et al., 2024), which is a benchmark used to evaluate code generation performance on pragmatic code generation tasks, i.e., code generation with repository context. It consists of 230 Python and 230 Java tasks from real-world open-source projects. Each task contains a function signature [^2] We use “method” and “function” interchangeably in this paper., a task description, a solution code as the ground truth, and several unit tests to assess the functional correctness of the generated code. It is worth noting that the repositories to which these Python tasks belong do not have specific requirements for code style. The objective of each task is to complete the code specified by the function signature, guided by the provided task description, and ensure that it passes the associated unit tests. In this study, we focus on Python tasks due to Python’s popularity (Srinath, 2017) and its alignment with previous code generation work (Bafatakis et al., 2019).
3.3. Implementation Details
During the inference stage, for the four Code LLMs, we employ a random sampling strategy, setting the maximum context length to 1,024 and the temperature to 0.8. For GPT-4, we adopted identical hyperparameter settings. The choices for maximum context length and temperature are aligned with the experimental setup used in CoderEval (Yu et al., 2024). All experiments are conducted on a machine with 216 GB of RAM and a Tesla A100 GPU with 80 GB of memory.